OpenStudy (ilovepurple42043):

the table of values below shows the domain and range of which function?

OpenStudy (ilovepurple42043):



OpenStudy (ilovepurple42043):

A. f(x)=a^x B. f(x)=alogx+b C. f(x)=\[\sqrt{x+b+c}\] D.f(x)=ax^2+bx+c